# The Early Bird Catches the Leak: Unveiling Timing Side Channels in LLM Serving Systems

### v4

@@ Abstract
Index Terms:
LLM, KV cache, Semantic cache, Side channels
@@ I Introduction
Large Language Models (LLMs) are widely used in applications such as chatbots [1, 2], search engines [3], and coding assistants [4]. However, LLM inference is resource-intensive, requiring substantial computational power and memory due to the model’s vast parameters, numerous layers, and large context sizes. Improving LLM inference performance has thus become essential, leading to solutions such as weight quantization [5, 6, 7, 8, 9], model compression [10, 11, 12], algorithm optimization [13, 14, 15], hardware advancements [16, 17], and parallel processing techniques [18]. These approaches aim to reduce latency and improve inference efficiency [19], though their privacy implications remain less clear.
@@
LLM cache channels. In our work, we examined various caches in LLM systems, which not only reduce the computational cost of LLM inference but also improve user experience by lowering service latency. We found that these caches can be misused to infer proprietary system prompts or sensitive prompts from peer users. These prompts may contain private user information and also hold commercial value, as they enable an LLM to carry out various downstream tasks without additional fine-tuning. We identified two primary cache channels:
$\bullet$ Leakage from the KV cache. For each inference request, the LLM maintains an in-memory state called the KV cache, which is reused in every iteration throughout the request’s entire service time. Due to the causal attention mask in LLMs, each token’s activations are influenced only by preceding tokens in the sequence. Thus, if multiple requests share a common prefix, the key and value embeddings for those prefix tokens are identical across sequences. To optimize the KV cache’s memory usage, the system identifies matching prompt prefixes across multiple requests and shares their key and value embeddings in memory at runtime [14, 20]. This sharing occurs when prompts include a common prefix, which frequently happens with few-shot examples [21], chatbot system prompts [22], or prompt templates [23]. For example, it has been noted that Claude’s prompt caching feature can reduce costs by up to 90% and decrease latency by up to 85% for long prompts [24].
@@
To address these challenges, we developed various attack strategies to exploit LLM side channels. Specifically, we use a threshold-based classification model to detect token-level KV cache hits based on offline timing measurements. We observe that online detection accuracy can be substantially improved with only a few repeated trials. To reduce the search space for the KV cache channel, we propose an incremental search algorithm that capitalizes on the requirement for prompts to share a common prefix, allowing us to recover the victim’s prompt token by token. For the semantic cache channel, we design an algorithm to select the most representative prompts as the attacker’s requests, given a targeted semantic focus. To minimize interference from the attacker’s own requests, we introduce a mechanism to clear cached data via batches of irrelevant requests. For the semantic cache, our method also ensures the attacker’s requests remain distinct by computing their semantic similarities.
@@
Experimental studies. In our study, we verified the presence of timing leakages in open-source projects, including SGLang [26], Langchain [27], and GPTCache [28]. Building on these findings, we demonstrate the feasibility of deducing proprietary system prompts (i.e., prompt stealing attack) and inferring sensitive requests from neighboring users (i.e., peeping neighbor attack).
@@
For the prompt stealing attack, our evaluation indicates that the accuracy of detecting per-token cache hits or misses in the KV cache is 99%, with a false positive rate (FPR) of 0.003. Using the incremental search algorithm, we recovered the system prompt token by token, requiring an average of 111.46 queries per recovered token. This approach achieved an average recovery accuracy of 89.0% and a corresponding FPR of 0.04. For the peeping neighbor attack, our measurements show an 81.4% accuracy in distinguishing hits from misses, with an average FPR of 0.045 in a single trial. This accuracy improved to 95.4% with a 0.056 FPR after 5 trials under GPTCache’s default settings. We further observed that it is possible to infer the documents processed by a victim user in a vulnerable LLM application, even when using standard commodity LLM services. Moreover, our black-box study of existing online services shows that popular LLM systems—such as Claude, DeepSeek, and Azure OpenAI—employ KV or semantic cache sharing to cut costs, rendering them susceptible to timing side-channel attacks.
@@
Finally, we propose initial defenses against these side-channel risks. To mitigate KV cache leakage, we recommend sharing prefix caches only in batches of at least $k$ tokens ($k=2,3,4$, etc.). Although this increases the prompt search space and thus the required number of guesses, the larger timing differences for sharing multiple tokens also make classifiers more robust. Consequently, attacks remain accurate but incur higher query overhead. To address semantic cache leakage, we advise anonymizing privacy-related content in user inputs before performing semantic-similarity searches. Preliminary experiments show that this measure adds modest overhead (around 4%).
@@
$\bullet$ New discovery. We identified new timing side channels in both open-source and online LLM serving systems, arising from the sharing of KV caches and semantic caches to lower inference costs.
@@
Responsible disclosure. We disclosed our findings to all relevant developers (SGLang, GPTCache, etc.) and LLM service providers (OpenAI, Claude, Google Gemini, etc.) upon identifying the side channels in September 2024. At the time of this manuscript’s preparation, we received positive responses from the SGLang team, which noted that we were among the first two groups to report this issue, both within the same week. Moreover, we were the first to raise the topic during the SGLang development meeting, and we are now working closely with their team on a resolution.
@@
Inspired by our findings, Gu et al. conducted a subsequent large-scale measurement study on prompt caching in real-world LLM services [31], identifying its presence in 8 out of 17 evaluated providers.
More recently, Zheng et al. investigated timing side channels in LLMs through the InputSnatch attack [30]. However, their work lacks the optimized search strategy we introduce for efficient request recovery and suffers from practical limitations. Specifically, their KV cache attacks on vLLM can extract blocks of 16 tokens but are limited to template-dependent scenarios, requiring attackers to have prior knowledge of fixed input structures. Additionally, their semantic cache attacks on GPTCache aim to infer patterns from retrieved documents but do not address self-induced interference from the attackers’ own queries and lack eviction strategies to mitigate adversarial noise. In contrast, our approach enables template-free extraction, independent of application-specific formats, and incorporates eviction-controlled probing to systematically eliminate interference. Furthermore, we validate the effectiveness of our method in a realistic setting: a document summarization service powered by a commercial LLM API.
@@ II-A LLM Serving Systems
In this paper, we explore the deployment of a shared LLM to serve multiple users or applications within a computing system. This setup is frequently observed in public services offered by commercial companies (e.g., OpenAI’s ChatGPT) and also applies to locally deployed shared enterprise LLMs, which are tailored to handle specific tasks, process large volumes of proprietary data, and meet unique business requirements. Additionally, the rise of LLM-based applications—often referred to as AI agents or co-pilots—has introduced a novel software paradigm that merges the capabilities of LLMs with traditional software functionalities. With the emergence of LLMs as operating systems (e.g., AIOS [32]) and agents functioning as apps, multiple LLM-based applications or agents can operate on the same shared LLM, treating it as a foundational model. These LLM agents are typically developed and deployed by different teams or organizations. This concept also extends to local LLM instances in a browser environment. For example, Lumos [33] is a Chrome extension powered by Ollama, a Retrieval-Augmented Generation (RAG) LLM co-pilot for web browsing, running entirely on local hardware without relying on remote servers.
In these scenarios, LLMs are typically optimized to achieve efficient latency and throughput, focusing on memory usage optimization, effective batching, and scheduling. However, complications arise from memory sharing, cache contention and eviction, and GPU scheduling across different users and applications. Such factors can introduce interference among concurrent requests, potentially leading to observable timing side channels. As these users and applications are not all mutually trusted, sensitive information leakage becomes a concern. This includes the potential exposure of other users’ confidential data—such as sensitive queries, proprietary system prompts, and processed documents—through timing side channels.
II-B Serving Frontend
@@
LLM serving modes. The LLM service offers two operation modes. In non-streaming mode, the response is fully generated and then delivered once the request has been processed. However, for long completions, this approach can result in an extended waiting period, possibly lasting several seconds. To achieve faster responses, the streaming mode is available. In this mode, the LLM emits tokens sequentially, allowing users to view the beginning of the completion while the remaining tokens are still being generated. Streaming is the preferred method for interacting with LLMs, especially in chatbot scenarios where real-time conversation is essential. Popular LLM applications (e.g., Bing Copilot [34], ChatGPT [1]) use a system prompt containing task definitions, examples, and safety rules to guide their behavior. This prompt is typically static and shared among all users.
@@
Metrics. Latency measures how long it takes for an LLM to respond to a user’s query, shaping users’ perceptions of speed and efficiency in generative AI applications. Low latency is particularly important for real-time interactions, such as chatbots and AI copilots. Time to First Token (TTFT) is the interval from the moment a user submits a prompt until receiving the first token of the response. It reflects the initial processing delay and serves as a crucial indicator of user-perceived responsiveness. Throughput, on the other hand, represents how many requests or tokens an LLM can process within a given time window. Since requests per second is affected by the model’s total generation time—which depends on output length—tokens per second is often used as the key metric for measuring throughput. This paper examines the risks arising from optimizing an LLM’s serving latency and employs TTFT as the primary metric for side-channel observations.
II-C Serving Backend
@@
Most LLMs rely on the Transformer architecture, which uses the attention mechanism [35] to pinpoint the most relevant parts of the input. Core to this mechanism are Query (Q), Key (K), and Value (V) embeddings: Q represents what the model is seeking at the current position, K encodes how to match relevant information across the sequence, and V holds the actual data to be retrieved when a match occurs. Leveraging scaled dot-product attention, the model processes Q, K, and V to selectively focus on the most pertinent parts of the input. LLM inference consists of two stages: the prefill phase and the decoding phase. The prefill phase processes the entire request prompt to produce the first output token, while the decoding phase generates subsequent tokens one by one.
Prefill phase. During the prefill phase, the LLM takes the request prompt as input and converts it into a sequence of tokens. Each token is transformed into a numerical representation, called an embedding, which the model can process. In this phase, the LLM computes the K and V embeddings for each token across every attention layer, enabling the generation of the first token of the response in a single step.
Decoding phase. In the decoding phase, the LLM generates each subsequent token by using the prefilled information and the single token produced in the previous step. For every layer, the engine computes the Q, K, and V embeddings for the new token and performs attention against all existing context tokens. Unlike the prefill phase, the decoding phase processes only one token at a time.
Memory management of KV cache. The attention mechanism in LLMs requires computing pairwise similarities among tokens in an input sequence, which leads to quadratic complexity with respect to sequence length [36]. To address this, KV caching stores the key and value embeddings in GPU memory, eliminating redundant computations and allowing the computation cost to scale linearly with sequence length.
Originally, LLM serving systems would statically allocate a sizable portion of memory for storing the KV cache, due to the unpredictable lengths of model outputs. However, this led to significant internal and external fragmentation. To mitigate these issues, vLLM introduced PagedAttention, which divides the KV cache into blocks and accesses them through a lookup table [14]. This table maps virtual cache blocks to physical locations in GPU memory, enabling efficient memory sharing across different requests. Modern LLM inference frameworks such as Nvidia’s TensorRT-LLM [37] and Huggingface’s TGI [38] incorporate similar concepts, but the security implications of sharing KV caches have not been thoroughly studied, leaving a critical gap in existing research.
II-D Threat Model
@@
In this paper, we examine the security implications of deploying a shared LLM to serve multiple users or applications within a single computing system. Specifically, we consider two main scenarios. First, an LLM service provider offers public APIs that registered users can employ to send requests, all of which are processed by the same underlying serving system. In this context, a victim user may establish a proprietary system prompt to power a widely used LLM application. Meanwhile, an attacker could leverage the same LLM APIs to infer this system prompt, thereby gaining potential financial benefits or circumventing the safety instructions encoded in the prompt. Second, public LLM applications—such as chatbots (e.g., OpenAI’s GPT-4) or document analysis services (e.g., AnythingLLM [39], Klu [40], etc.)—handle concurrent requests from multiple users via the same LLM serving system. If the application itself relies on a public LLM API, these requests are generally routed through the same developer’s API key. An attacker could register as a user of such an application to discover whether specific requests have been submitted by others. For instance, they might seek to determine whether a user has shown interest in a particular topic or uploaded a specific file. They could also monitor requests over time to detect private attributes or sensitive personally identifiable information (PII) (Table III). In both scenarios, the attacker’s and the victim’s requests share the same platform and thus make use of the same caches. This shared environment can produce interference and create observable timing side channels—the core subject of our investigation. The attacker needs only black-box access to the underlying model, without knowledge of its architecture or weight parameters. However, the attacker must first examine the system’s leakage profile in an offline phase, analyzing how different inputs affect timing. This analysis helps them craft inputs that exploit the timing discrepancies introduced by cache sharing.
@@
In this paper, we explore the side-channel risks linked to both local and remote LLM services. For local settings, we assume a stable network connection between client and server. For remote services, previous works—such as NetCAT [41] and NetSpectre [42]—have addressed mitigating noise caused by unstable connections and jitters, particularly in CPU cache side-channel attacks. Extending such noise-reduction strategies to remote LLM scenarios remains an avenue for future research. We do not consider hardware side channels tied to GPU micro-architectures (e.g., GPU caches [43], residue-based leakage [44], or power/frequency side channels [45]). Instead, our focus lies on software caches maintained by the LLM serving system, making our attacks applicable across various hardware platforms (CPUs, GPUs, ASICs, etc.).

### v5

@@ I Introduction
Large Language Models (LLMs) are widely used in applications such as chatbots [1], search engines [2], and coding assistants [3]. However, LLM inference is resource-intensive, requiring substantial computational power and memory due to the model’s vast parameters, numerous layers, and large context sizes. Improving LLM inference performance has thus become essential, leading to solutions such as weight quantization [4, 5, 6], model compression [7], and algorithm optimization [8, 9]. While these approaches reduce latency and improve inference efficiency, their privacy implications remain less clear.
@@
LLM cache channels. We examined various caches in LLM systems, which not only reduce the computational cost of LLM inference but also improve user experience by lowering service latency. We found that these caches can be misused to infer proprietary system prompts or sensitive prompts from peer users. These prompts may contain private user information and also hold commercial value, as they enable an LLM to carry out various downstream tasks without additional fine-tuning. We identified two primary cache channels:
$\bullet$ Leakage from the KV cache. For each inference request, the LLM maintains an in-memory state called the KV cache, which is reused throughout the request’s entire service time. Due to the causal attention mask in LLMs, each token’s activations are influenced only by preceding tokens in the sequence. Thus, if multiple requests share a common prefix, the key and value embeddings for those prefix tokens are identical across sequences. To optimize the KV cache’s memory usage, the system identifies matching prompt prefixes across multiple requests and shares their key and value embeddings in memory at runtime [9, 10]. This sharing occurs when prompts include a common prefix, which frequently happens with few-shot examples, chatbot system prompts, or prompt templates [11]. For example, it has been noted that Claude’s prompt caching feature can reduce costs by up to 90% and decrease latency by up to 85% for long prompts [12].
@@
To address these challenges, we developed various attack strategies to exploit LLM side channels. Specifically, we use a threshold-based classification model to detect token-level KV cache hits based on offline timing measurements. We observe that online detection accuracy can be substantially improved with only a few repeated trials. To reduce the search space for the KV cache channel, we propose an incremental search algorithm that capitalizes on the requirement for prompts to share a common prefix, allowing us to recover the victim’s prompt token by token. For the semantic cache channel, we design an algorithm to select the most representative prompts as the attacker’s requests, given a targeted semantic focus. To minimize interference from the attacker’s own requests, we introduce a mechanism to clear cached data via batches of irrelevant requests. Our method also ensures the attacker’s requests remain distinct by computing their semantic similarities.
@@
Experimental studies. In our study, we confirmed timing leakages in open-source projects, including SGLang [10], Langchain [14], and GPTCache [13]. We further demonstrate the feasibility of deducing proprietary system prompts (i.e., prompt stealing attack) and inferring sensitive requests from neighboring users (i.e., peeping neighbor attack).
@@
For the prompt stealing attack, our evaluation indicates that the accuracy of detecting per-token cache hits or misses in the KV cache is 99%, with a false positive rate (FPR) of 0.003. Using the incremental search algorithm, we recovered the system prompt token by token, requiring an average of 112.72 queries per recovered token. This approach achieved an average recovery accuracy of 89.0% and a corresponding FPR of 0.04. For the peeping neighbor attack, our measurements show an 81.4% accuracy in distinguishing hits from misses, with an average FPR of 0.045 in a single trial. This accuracy improved to 95.4% with a 0.056 FPR after 5 trials under GPTCache’s default settings. We further observed that it is possible to infer the documents processed by a victim user in a vulnerable LLM application, even when using standard commodity LLM services. Moreover, our black-box study of existing online services shows that popular LLM systems—such as Claude, DeepSeek, and Azure OpenAI—employ KV or semantic cache sharing to cut costs, rendering them susceptible to timing side-channel attacks.
@@
Finally, we propose initial defenses against these side-channel risks. To mitigate KV cache leakage, we recommend sharing prefix caches only in batches of at least $k$ tokens ($k\geq 2$). Although this increases the prompt search space and thus the required number of guesses, the larger timing differences for sharing multiple tokens also make classifiers more robust. Consequently, attacks remain accurate but incur higher query overhead. To address semantic cache leakage, we advise anonymizing privacy-related content in user inputs before performing semantic-similarity searches. Our experiments show that this method adds modest overhead (around 4%).
@@
$\bullet$ New discovery. We identified new timing side channels in both open-source and online LLM serving systems, arising from the sharing of KV caches and semantic caches.
@@
Responsible disclosure. We disclosed our findings to all relevant developers (SGLang, GPTCache, etc.) and LLM service providers (OpenAI, Claude, etc.) upon identifying the side channels in September 2024. At the time of this manuscript’s preparation, we received positive responses from the SGLang team, which noted that we were among the first two groups to report this issue, both within the same week. Moreover, we were the first to raise the topic during the SGLang development meeting, and we are now working closely with their team on a resolution.
@@
Inspired by our findings, Gu et al. conducted a subsequent large-scale measurement study on prompt caching in real-world LLM services [17], identifying its presence in 8 out of 17 evaluated providers. More recently, Zheng et al. investigated timing side channels in LLMs through the InputSnatch attack [16]. However, their work lacks the optimized search strategy we introduce for efficient request recovery and suffers from practical limitations. Specifically, their KV cache attacks on vLLM can extract blocks of 16 tokens but are limited to template-dependent scenarios, requiring attackers to have prior knowledge of fixed input structures. Additionally, their semantic cache attacks on GPTCache aim to infer patterns from retrieved documents but do not address self-induced interference from the attackers’ own queries and lack eviction strategies to mitigate adversarial noise. In contrast, our approach enables template-free extraction, independent of application-specific formats, and incorporates eviction-controlled probing to systematically eliminate interference. Furthermore, we validate the effectiveness of our method in a realistic setting: a document summarization service powered by a commercial LLM API.
In general, PROMPTLEAK and InputSnatch achieve higher prompt recovery rates and lower per-token recovery costs than our attack, but at the expense of much stronger assumptions. Specifically, PROMPTLEAK depends on the LPM scheduling policy in SGLang, while InputSnatch relies on predefined template structures and fixed sentence blocks known to the attacker. By contrast, our method only assumes cache sharing and does not require such constraints. Even under this weaker and more realistic model, our attack achieves a reasonable single-token recovery accuracy and can recover a meaningful number of tokens — 81 tokens with only 519 guesses (Section IV-A). In practice, attackers could recover even more tokens by employing a stronger next-token predictor and increasing the maximum number of attack queries.
@@ II-A LLM Serving Systems
In this paper, we explore shared LLM deployments serving multiple users or applications within computing systems. This setup is common in commercial public services (e.g., ChatGPT) and local LLMs tailored to specific tasks. In these scenarios, LLMs optimize latency and throughput via memory usage optimization, efficient batching, and GPU scheduling. These factors can introduce interference among concurrent requests, potentially leading to observable timing side channels.
@@
LLM serving modes. The LLM service offers two operation modes. In non-streaming mode, the response is fully generated and then delivered once the request has been processed. However, this approach can result in an extended waiting period for long completions. To achieve faster responses, the streaming mode is available. In this mode, the LLM emits tokens sequentially, allowing users to view the beginning of the completion while the remaining tokens are still being generated. Streaming is the preferred method for interacting with LLMs, especially in chatbot scenarios where real-time conversation is essential. Popular LLM applications (e.g., ChatGPT) use a system prompt containing task definitions, examples, and safety rules to guide their behavior. This prompt is typically static and shared among all users.
@@
Metrics. Latency measures how long it takes for an LLM to respond to a user’s query. Low latency is particularly important for real-time interactions, such as chatbots. Time to First Token (TTFT) is the interval from the moment a user submits a prompt until receiving the first token of the response. It reflects the initial processing delay and serves as a crucial indicator of user-perceived responsiveness. This paper examines the risks arising from optimizing an LLM’s serving latency and employs TTFT as the primary metric for side-channel observations.
II-B Serving Backend
@@
Most LLMs rely on the Transformer architecture, which uses Attention [18] to pinpoint the most relevant parts of the input. Core to this mechanism are Query (Q), Key (K), and Value (V) embeddings: Q specifies what to seek at each position, K determines how to match relevant information across the sequence, and V provides the data retrieved upon a match. LLM inference consists of two stages: the prefill phase and the decoding phase. In the prefill phase, the LLM processes the request prompt by converting it into a sequence of tokens, each mapped to a numerical representation called embedding. It then computes the K and V embeddings for each token across all attention layers, enabling the generation of the first token in a single step. In the decoding phase, the LLM generates each subsequent token by using the prefilled context and the previously generated token. For each layer, it computes the Q, K and V embeddings for the new token and attends to all existing context tokens. Unlike the prefill phase, decoding processes only one token at a time.
Memory management of KV cache. The attention mechanism in LLMs requires computing pairwise similarities among tokens, which leads to quadratic complexity with respect to sequence length. To address this, KV caching stores the K and V embeddings in GPU memory, avoiding redundant computations and reducing the computation cost to scale linearly with sequence length. Early LLM systems statically allocated large memory regions for KV caches to accommodate unpredictable output lengths, leading to significant internal and external fragmentation. vLLM mitigated this with PagedAttention, which divides the KV cache into blocks accessed via a lookup table [9]. This enables efficient memory sharing across requests. Modern frameworks like Nvidia’s TensorRT-LLM and Huggingface’s TGI adopt similar techniques, but the security implications of KV cache sharing remain underexplored, leaving a critical gap in existing research.
II-C Threat Model and Assumptions
@@
Threat model. In this paper, we examine the security risks of deploying a shared LLM to serve multiple users or applications within a single computing system. Specifically, we consider two main scenarios. First, an LLM service provider offers public APIs that registered users can employ to send requests, all of which are processed by the same underlying serving system. A victim user may establish a proprietary system prompt to power a popular LLM application, while an attacker could leverage the same LLM APIs to infer this system prompt, thereby gaining financial benefits or bypassing embedded safety instructions. Second, public LLM applications—such as chatbots (e.g., GPT-4) or document analysis services (e.g., AnythingLLM [19], etc.)—handle concurrent user requests via the same LLM serving system. If the application itself relies on a public LLM API, these requests are generally routed through the same developer’s API key. An attacker could register as a user of the application to detect whether specific requests have been submitted by others and infer, for instance, users’ interest in certain topics or uploaded files. They could also monitor requests over time to detect private attributes or sensitive personally identifiable information (PII) (Table II). In both scenarios, the attacker’s and the victim’s requests share the same platform and thus same cache resources. This shared environment can produce interference and create observable timing side channels—the core subject of our investigation. The attacker needs only black-box access to the underlying model, without knowledge of its architecture or weight parameters. However, the attacker must first examine the system’s leakage profile in an offline phase to understand how different inputs affect timing, enabling them to craft queries that exploit the cache-induced timing discrepancies.
@@
In this paper, we explore the side-channel risks in both local and remote LLM services. For local settings, we assume a stable network connection between client and server. For remote services, previous works such as NetCAT [20] and NetSpectre [21] have addressed mitigating noise caused by unstable connections and jitters, particularly in CPU cache side-channel attacks. Extending such noise-reduction strategies to remote LLM scenarios remains an avenue for future research. We do not consider hardware side channels tied to GPU micro-architectures [22, 23, 24]. Instead, our focus lies on software caches maintained by the LLM serving system, making our attacks applicable across various hardware platforms (CPUs, GPUs, ASICs, etc.).
Assumptions. For end-to-end attacks, we assume that the adversary has certain prior knowledge of the target user or system prompt. Specifically:
  • Prompt Stealing Attacks (Section IV-A): The attacker is assumed to have access to a public dataset of system prompts that can be used to fine-tune the next-token predictor. Such datasets are semantically diverse and representative of real-world prompt engineering practices. In our evaluation, the victim randomly selects a small subset (e.g., 200 prompts) as its proprietary system prompt, while the attacker uses the remaining prompts as training data. This reflects realistic scenarios where production systems employ a limited number of carefully crafted prompts, while public repositories offer abundant similar data for adversarial training.
  • Peeping Neighbor Attacks (Section IV-B): We assume the attacker is focused on a particular topic (e.g., healthcare, travel planning) and has a public set of prompt templates containing sensitive private attributes such as personal names, medical conditions, travel destinations, or business details. These templates can be obtained from public datasets or generated using paraphrasing tools. The attacker uses them to probe semantic caches, where timing differences between cache hits and misses may reveal whether the victim’s request contains the targeted attributes.
  • Document Inference Attacks (Section IV-C): Because KV-cache sharing typically requires requests to share a common prefix, we assume the attacker knows either the full document or at least its prefix (i.e., the initial portion of the content). Even if the document’s content is known, detecting its submission to an LLM service may disclose an organization’s interests, posing serious privacy and competitive risks across different sectors.